# Stability of Non-Hyperbolic Sets

### v1

@@ Acknowledgements.
The authors express their gratitude to Michael Benedicks for his continued support throughout the preparation of this paper and for the many invaluable discussions at both the inspirational and technical levels. The second author was partially supported by the starting grant SFB 67575 and by the VR grant 67578, Renormalization in dynamics.
@@ 2 Preliminaries
$\lambda\neq\mu^{k_{1}}\text{ and }\mu\neq\lambda^{k_{2}}$ (2.1)
@@
Definition 1.
Let $\mathcal{M}$ be an two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a diffeomorphism with saddle point $p\in\mathcal{M}$. We say that $p$ satisfies the ${\mathcal{C}}^{4}$} non-resonance condition if (2.1) holds.
Theorem 2.
Let $\mathcal{M}$ be a two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a diffeomorphism with a saddle point $p\in\mathcal{M}$ which satisfies the ${\mathcal{C}}^{4}$} non-resonance condition. Let $0\in\mathcal{P}\subset\mathbb{R}^{n}$ and let $F:\mathcal{P}\times\mathcal{M}\to\mathcal{M}$ be a $\mathcal{C}$$^{\infty}$} family with $F_{0}=f$. Then, there exists a neighborhood $U$ of $p$ and a neighborhood $V$ of 0 such that, for every $t\in V$, $F_{t}$ has a saddle point $p_{t}\in U$ satisfying the ${\mathcal{C}}^{4}$} non-resonance condition. Moreover $p_{t}$ is ${\mathcal{C}}^{4}$} linearizable in the neighborhood $U$ and the linearization depends ${\mathcal{C}}^{4}$} on the parameters.
The proofs of Theorem 1 and Theorem 2 can be found in [12, 26].
2.1 Families of unfoldings
In the sequel we define a map with a strong homoclinic tangency by listing several conditions. To summarize, a map with a strong homoclinic tangency has a non degenerate homoclinic tangency $q_{1}$ and a transversal homoclinic intersection $q_{2}$. Conditions $(f6)$, $(f7)$ and $(f8)$ ensure that $q_{1}$ and the unstable local manifolds of $q_{1}$ and $q_{2}$ accumulate on the leg of the unstable manifold of the saddle point containing the transversal homoclinic intersection $q_{2}$. These conditions are redundant if $\mu<-1$. Moreover to implement the construction, a condition on the eigenvalues is also required, see $(f2)$. The basic properties of a map with a strong homoclinic tangency are summarized in Figure 1.
Definition 2.
@@ 2.1 Families of unfoldings
Remark 2.
Observe that all conditions are open in the space of maps with an homoclinic and transversal tangency. Also, except for $(f2)$, all conditions are dense.
Remark 3.
@@
Following [35], we construct now a family of unfolding of a map $f$ with a homoclinic tangency and a transversal intersection. Let $\mathcal{P}=[-r,r]\times[-r,r]$ with $r>0$. Given a map $f$ with a strong homoclinic tangency, we consider a $\mathcal{C}$$^{\infty}$} family $F:\mathcal{P}\times\mathcal{M}\to\mathcal{M}$ through $f$ with the following properties:
@@
    $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1,$ (2.2)
@@
$0<\frac{\log\mu_{\text{max}}}{\log{\frac{1}{\lambda_{\text{max}}}}}<\frac{3}{2}\frac{\log\mu_{\text{min}}}{\log{\frac{1}{\lambda_{\text{min}}}}}<\frac{1}{2}.$ (2.3)
@@ 2.1.1 Rescaled Families
According to Theorem 2 we may make a change of coordinates to ensure that the family $F$ is ${\mathcal{C}}^{4}$} and for all $(t,a)\in[-r_{0},r_{0}]\times[-r_{0},r_{0}]$ with $0<r_{0}<r$ and by rescaling we can assume that $F_{t,a}$ is linear on the ball $2,2]^{2}$, namely
$F_{t,a}=\left(\begin{matrix}\lambda(t,a)&0\\ 0&\mu(t,a)\\ \end{matrix}\right).$ (2.4)
@@ 2.1.2 Critical Points and Critical Values
The following lemma says that $F^{N}_{t,a}(x,y)$ produces an unfolding in the sense of [35]. The formal definition of unfolding is given in Definition 4 below.
Lemma 1.
@@
$\frac{\partial Y_{t,a}}{\partial y}\left(x,c(x,t,a)\right)=0,$ (2.5)
@@
Moreover
$|c(x,t,a)-c(0,t,a)|=O\left(|x|\right).$ (2.6)
@@
Another instance where critical points occur but play only a secondary role is in the context of strongly dissipative Hénon maps at the boundary of chaos. These maps have a period doubling Cantor attractor, see [14]. The Cantor attractors are studied using renormalization zooming in to the so-called tip of the Cantor set as in [14]. Indeed, return maps to neighborhoods of the tip are Hénon-like maps. A posterior one shows that the stable manifold at the tip is tangent to the direction of the neutral Lyapunov exponent. The tip plays the role of a critical point. However, this fact did not play any role during the renormalization analysis. This is another instance where, from a technical point of view, the notion of critical point is not that important.
Definition 4.
A family $F_{t,a}$ is called an unfolding of $f$ if it can be reparametrized such that
@@
Remark 6.
Without lose of generality by a suitable coordinate change we may assume that if $F$ is an unfolding then $z_{y}(t,a)=a$, the primary critical value is at height $a$ and the primary critical point $c(t,a)=(0,1)$, see Figure 2.
Remark 7.
A generic two-dimensional family trough f can locally be reparametrized to become an unfolding.
Remark 8.
@@ 2.2 Properties of unfoldings
$W_{n}(x)=\mu^{-n}W\left(\lambda^{n}x\right),$ (2.7)
@@
$\frac{1}{2\mu}\mu^{-n}\leq|W_{n}|\leq 2\mu^{-n},$ (2.8)
$\left|\frac{\partial W_{n}}{\partial x}\right|=O\left(\frac{\lambda}{\mu}\right)^{n},$ (2.9)
@@
$\left|\frac{\partial^{2}W_{n}}{\partial x^{2}}\right|=O\left(\frac{\lambda^{2}}{\mu}\right)^{n}.$ (2.10)
@@
Fix a point $x\in[-2,2]$ and a parameter $(t,a)\in[-r_{0},r_{0}]^{2}$. We denote by $W_{n}$ the manifold corresponding to $(t,a)$ and by $W_{n}+\Delta W_{n}$ the manifold to $(t,a+\Delta a)$. Then, since the maps $F^{n}_{t,a}$ are linear, by differentiating (2.7) we obtain
@@
$\frac{da_{n}}{dt}=-n\frac{\partial\mu}{\partial t}\frac{1}{\mu^{n+1}}\left[1+O\left(\lambda^{n}\right)\right].$ (2.11)
@@
Then $\Gamma$ is a ${\mathcal{C}}^{2}$} codimension $1$ manifold transversal to $W^{u}(0)$. For $n\geq 0$ let
$\Gamma_{n}=\left\{\left(x,y,t,a\right)|F^{n}_{t,a}\left(x,y\right)\in\Gamma\right\},$ (2.12)
@@
$\frac{\partial z_{y}}{\partial a}=1\text{ and }z_{y}(t,0)=0.$ (2.13)
Let $Z=\text{Image}(z)$. Because of (2.13), $Z$ is a manifold transversal to $\Gamma_{\infty}$. Hence, there exists $n_{1}>0$ such that, for all $n\geq n_{1}$, $Z$ is transversal to $\Gamma_{n}$. As consequence, for all $n\geq n_{1}$,
@@
is a ${\mathcal{C}}^{2}$} codimension $1$ manifold. We define $a_{n}:t\mapsto a_{n}(t)$ as a function whose graph is $A_{n}$. Observe that, by Lemma 1 and Remark 5, the ${\mathcal{C}}^{2}$} function $c$ satisfies, $c(0,t,a)=1$. Hence $\partial c/\partial t=O(x)$ and $\partial c/\partial a=O(x)$. Recall that
$\mu^{n}a_{n}=c\left(\lambda^{n}z_{x},t,a_{n}\right).$ (2.14)
By differentiating (2.14) and using $\partial c/\partial a,\partial c/\partial t=O(x)$ we have
@@
A proof similar to that of Lemma 2.7 gives the following.
Lemma 4.
Let $\Gamma_{n}$ be as in (2.12), then
@@
Exact form of the map $F^{N}$.
For families of unfoldings the following holds.
Lemma 5.
There exist $x^{\prime}_{0}<x_{0}$, $a^{\prime}_{0}<a_{0}$, $b>0$ and $Q>0$, such that, for all $(t,a)\in[-t_{0},t_{0}]\times[-a^{\prime}_{0},a^{\prime}_{0}]$ and for every $(x,y)\in\Gamma_{t,a}$ with $|x|<|x^{\prime}_{0}|$ the following holds. There exist matrices $A_{x,y}$, $B_{x,y}\neq 0$ and $C_{x,y}$ such that $F^{N}_{t,a}$ in coordinates centered in $(x,y)$ and $F^{N}_{t,a}(x,y)$ has the form
$F^{N}_{t,a}\left(\begin{matrix}\Delta x\\ \Delta y\end{matrix}\right)$ $=$ $\left(\begin{matrix}A_{x,y}&B_{x,y}\\ C_{x,y}&0\end{matrix}\right)\left(\begin{matrix}\Delta x\\ \Delta y\end{matrix}\right)$ (2.15)
@@
where the matrices $A_{x,y}$, $C_{x,y}$ and the vector $B_{x,y}$ are ${\mathcal{C}}^{2}$} dependent on $x$ and $y$, $Q_{x,y}>Q$, $|B_{x,y}|>b>0$ and the vector valued functions $O_{i,j}$ are ${\mathcal{C}}^{2}$} dependent on $x$ and $y$ and they are uniformly bounded.
@@
The lemma gives the Taylor expansions of $F^{N}_{t,a}$ when $(x,y)\in\Gamma_{t,a}$. It follows immediately from Lemma 1 where we state that a vertical curve trough $(x,y)$ in $\Gamma_{t,a}$ is mapped to a curve with a non degenerate horizontal tangency. In particular $\partial Y_{t,a}/\partial y=0$ by (2.5) and the horizontal tangency,
@@
is not degenerate for all $(x,y)\in\Gamma_{t,a}$. This is a consequence of the following argument. Let $t\in[-t_{0},t_{0}]$. Because $F^{N}_{t,0}(q_{3}(t))=q_{1}(t)$ is a non degenerate homoclinic tangency, we know that the vector $B_{q_{3}(t)}\neq 0$ and $Q_{q_{3}(t)}>0$. By taking $|a|<|a^{\prime}_{0}|$, $|x|<|x^{\prime}_{0}|$ small enough, the lower bounds on $Q_{x,y}$ and $B_{x,y}$ follow. ∎
@@ 2.3 Choice of $\theta$
In this subsection, we introduce a small number $\theta$ that is going to play a fundamental role in the construction of our combinatorics.
@@
$1<\lambda_{\text{min}}^{2\theta}\mu_{\text{min}}^{3}\text{ and }\lambda_{\text{max}}^{3\theta}\mu^{4}_{\text{max}}<1.$ (2.16)
@@
$0<\theta_{0}=\frac{4}{3}\frac{\log\mu_{\text{max}}}{\log{\frac{1}{\lambda_{\text{max}}}}}<\theta<\frac{3}{2}\frac{\log\mu_{\text{min}}}{\log{\frac{1}{\lambda_{\text{min}}}}}=\theta_{1}<\frac{1}{2},$ (2.17)
where we used $(F3)$, the initial condition $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1$ and (2.3).
Moreover, by restricting the parameter domain, we choose $\theta$ such that
$\left(\frac{\theta}{\lambda_{\text{max}}}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu_{\text{max}}}\frac{\partial\mu}{\partial t}\right)>0.$ (2.18)

### v3

@@ Acknowledgements.
The authors express their gratitude to Michael Benedicks for his continued support throughout the preparation of this paper and for the many invaluable discussions at both the inspirational and technical levels. The authors are also grateful to Michael Lyubich for shedding light on aspects of our results that we had previously overlooked, and for his valuable insights connecting the different parts of this work. The second author was partially supported by the starting grant SFB 67575 and by the VR grant 67578, Renormalization in dynamics.
@@ 2 Preliminaries
$\lambda\neq\mu^{k_{1}}\text{ and }\mu\neq\lambda^{k_{2}}$ (2.0.2)
@@
Definition 2.0.3.
Let $\mathcal{M}$ be an two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a diffeomorphism with saddle point $p\in\mathcal{M}$. We say that $p$ satisfies the ${\mathcal{C}}^{4}$} non-resonance condition if (2.0.2) holds.
Theorem 2.0.4.
Let $\mathcal{M}$ be a two dimensional $\mathcal{C}$$^{\infty}$} manifold and let $f:\mathcal{M}\to\mathcal{M}$ be a diffeomorphism with a saddle point $p\in\mathcal{M}$ which satisfies the ${\mathcal{C}}^{4}$} non-resonance condition. Let $0\in\mathcal{P}\subset\mathbb{R}^{n}$ and let $F:\mathcal{P}\times\mathcal{M}\to\mathcal{M}$ be a $\mathcal{C}$$^{\infty}$} family with $F_{0}=f$. Then, there exists a neighborhood $U$ of $p$ and a neighborhood $V$ of $0\in\mathcal{P}$ such that, for every $t\in V$, $F_{t}$ has a saddle point $p_{t}\in U$ satisfying the ${\mathcal{C}}^{4}$} non-resonance condition. Moreover $p_{t}$ is ${\mathcal{C}}^{4}$} linearizable in the neighborhood $U$ and the linearization depends ${\mathcal{C}}^{4}$} on the parameters.
The proofs of Theorem 2.0.1 and Theorem 2.0.4 can be found in [14, 38].
2.1 Unfoldings
In the sequel we define a map with a strong homoclinic tangency by listing several conditions. To summarize, a map with a strong homoclinic tangency has a non degenerate homoclinic tangency $q_{1}$ and a transversal homoclinic intersection $q_{2}$. Conditions $(f6)$, $(f7)$ and $(f8)$ ensure that $q_{1}$ and the unstable local manifolds of $q_{1}$ and $q_{2}$ accumulate on the leg of the unstable manifold of the saddle point containing the transversal homoclinic intersection $q_{2}$. These conditions are redundant if $\mu<-1$. Moreover to implement the construction, a condition on the eigenvalues is also required, see $(f2)$. The basic properties of a map with a strong homoclinic tangency are illustrated in Figure 1.
Definition 2.1.1.
@@ 2.1 Unfoldings
Remark 2.1.3.
Observe that all conditions are open in the space of maps with an homoclinic and transversal tangency.
Remark 2.1.4.
@@
Following [50], we consider now an unfolding of a map $f$ with a homoclinic tangency and a transversal intersection. Let $\mathcal{P}=[-r,r]\times[-r,r]$ with $r>0$. Given a map $f$ with a strong homoclinic tangency, we consider a $\mathcal{C}$$^{\infty}$} family $F:\mathcal{P}\times\mathcal{M}\to\mathcal{M}$ through $f$ with the following properties:
@@
    $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1,$ (2.1.5)
@@
$0<\frac{\log\mu_{\text{max}}}{\log{\frac{1}{\lambda_{\text{max}}}}}<\frac{3}{2}\frac{\log\mu_{\text{min}}}{\log{\frac{1}{\lambda_{\text{min}}}}}<\frac{1}{2}.$ (2.1.7)
@@ 2.1.1 Rescaled Families
According to Theorem 2.0.4 we may make a change of coordinates to ensure that the family $F$ is ${\mathcal{C}}^{4}$} and for all $(t,a)\in[-t_{0},t_{0}]\times[-a_{0},a_{0}]$ with $0<t_{0},a_{0}<r$ and by rescaling we can assume that $F_{t,a}$ is linear on the ball $2,2]^{2}$, namely
$F_{t,a}=\left(\begin{matrix}\lambda(t,a)&0\\ 0&\mu(t,a)\\ \end{matrix}\right).$ (2.1.8)
@@ 2.1.2 Critical Points and Critical Values
The following lemma says that $F^{N}_{t,a}(x,y)$ produces an unfolding in the sense of [50]. The formal definition of unfoldings is given in Definition 2.1.14 below.
Lemma 2.1.9.
@@
$\frac{\partial Y_{t,a}}{\partial y}\left(x,c(x,t,a)\right)=0,$ (2.1.10)
@@
Moreover,
$|c(x,t,a)-c(0,t,a)|=O\left(|x|\right).$ (2.1.11)
@@
Another instance where critical points occur but play only a secondary role is in the context of strongly dissipative Hénon maps at the boundary of chaos. These maps have a period doubling Cantor attractor, see [17]. The Cantor attractors are studied using renormalization zooming in to the so-called tip of the Cantor set as in [17]. Indeed, return maps to neighborhoods of the tip are Hénon-like maps. A posterior one shows that the stable manifold at the tip is tangent to the direction of the neutral Lyapunov exponent. The tip plays the role of a critical point. However, this fact did not play any role during the renormalization analysis. This is another instance where, from a technical point of view, the notion of critical point is not that important. A similar definition of critical point in a more general setting is given in [22].
Definition 2.1.14.
A family $F_{t,a}$ is called an unfoldings of $f$ if it can be reparametrized such that
@@
Remark 2.1.15.
Without lose of generality by a suitable coordinate change we may assume that if $F$ is an unfolding then $z_{y}(t,a)=a$, the primary critical value is at height $a$, $z_{x}(t,a)=1$ and the primary critical point $c_{t,a}=(0,1)$, see Figure 2.
Remark 2.1.16.
A generic two-dimensional family trough f can locally be reparametrized to become an unfoldings.
Remark 2.1.17.
@@ 2.2 Properties of unfoldings
$W_{n}(x)=\mu^{-n}W\left(\lambda^{n}x\right),$ (2.2.1)
@@
$\frac{1}{2\mu}\mu^{-n}\leq|W_{n}|\leq 2\mu^{-n},$ (2.2.2)
$\left|\frac{\partial W_{n}}{\partial x}\right|=O\left(\frac{\lambda}{\mu}\right)^{n},$ (2.2.3)
@@
$\left|\frac{\partial^{2}W_{n}}{\partial x^{2}}\right|=O\left(\frac{\lambda^{2}}{\mu}\right)^{n}.$ (2.2.4)
@@
Fix a point $x\in[-2,2]$ and a parameter $(t,a)\in[-r_{0},r_{0}]^{2}$. We denote by $W_{n}$ the manifold corresponding to $(t,a)$ and by $W_{n}+\Delta W_{n}$ the manifold to $(t,a+\Delta a)$. Then, since the maps $F^{n}_{t,a}$ are linear, by differentiating (2.2.1) we obtain
@@
$\frac{da_{n}}{dt}=-n\frac{\partial\mu}{\partial t}\frac{1}{\mu^{n+1}}\left[1+O\left(\lambda^{n}\right)\right].$ (2.2.7)
@@
Then $\Gamma$ is a ${\mathcal{C}}^{2}$} codimension-11 manifold transversal to $W^{u}(0)$. For $n\geq 0$ let
$\Gamma_{n}=\left\{\left(x,y,t,a\right)|F^{n}_{t,a}\left(x,y\right)\in\Gamma\right\},$ (2.2.8)
@@
$\frac{\partial z_{y}}{\partial a}=1\text{ and }z_{y}(t,0)=0.$ (2.2.9)
Let $Z=\text{Image}(z)$. Because of (2.2.9), $Z$ is a manifold transversal to $\Gamma_{\infty}$. Hence, there exists $n_{1}>0$ such that, for all $n\geq n_{1}$, $Z$ is transversal to $\Gamma_{n}$. As consequence, for all $n\geq n_{1}$,
@@
is a ${\mathcal{C}}^{2}$} codimension-11 manifold. We define $a_{n}:t\mapsto a_{n}(t)$ as a function whose graph is $A_{n}$. Observe that, by Lemma 2.1.9 and Remark 2.1.12, the ${\mathcal{C}}^{2}$} function $c$ satisfies, $c(0,t,a)=1$. Hence $\partial c/\partial t=O(x)$ and $\partial c/\partial a=O(x)$. Recall that
$\mu^{n}a_{n}=c\left(\lambda^{n}z_{x},t,a_{n}\right).$ (2.2.10)
By differentiating (2.2.10) and using $\partial c/\partial a,\partial c/\partial t=O(x)$ we have
@@
A proof similar to that of Lemma 2.2.1 gives the following.
Lemma 2.2.11.
Let $\Gamma_{n}$ be as in (2.2.8), then
@@
Form of the map $F^{N}$.
For unfoldings the following holds.
Lemma 2.2.12.
There exist $x^{\prime}_{0}<x_{0}$, $a^{\prime}_{0}<a_{0}$, $b>0$ and $Q>0$, such that, for all $(t,a)\in[-t_{0},t_{0}]\times[-a^{\prime}_{0},a^{\prime}_{0}]$ and for every $(x,y)\in\Gamma_{t,a}$ with $|x|<|x^{\prime}_{0}|$ the following holds. There exist ${\mathcal{C}}^{3}$} functions $A_{x,y}$, $B_{x,y}\neq 0$ and $C_{x,y}$ such that $F^{N}_{t,a}$ in coordinates centered in $(x,y)$ and $F^{N}_{t,a}(x,y)$ has the form
$F^{N}_{t,a}\left(\begin{matrix}\Delta x\\ \Delta y\end{matrix}\right)$ $=$ $\left(\begin{matrix}A_{x,y}&B_{x,y}\\ C_{x,y}&0\end{matrix}\right)\left(\begin{matrix}\Delta x\\ \Delta y\end{matrix}\right)$ (2.2.13)
@@
where $Q_{x,y}>Q$, $|B_{x,y}|>b>0$ and the ${\mathcal{C}}^{2}$} functions $O_{i,j}$ are uniformly bounded.
@@
The lemma gives the Taylor expansions of $F^{N}_{t,a}$ when $(x,y)\in\Gamma_{t,a}$. It follows immediately from Lemma 2.1.9 where we state that a vertical curve trough $(x,y)$ in $\Gamma_{t,a}$ is mapped to a curve with a non degenerate horizontal tangency. In particular $\partial Y_{t,a}/\partial y=0$ by (2.1.10) and the horizontal tangency,
@@
is not degenerate for all $(x,y)\in\Gamma_{t,a}$. This is a consequence of the following argument. Let $t\in[-t_{0},t_{0}]$. Because $F^{N}_{t,0}(q_{3}(t))=q_{1}(t)$ is a non degenerate homoclinic tangency, we know that $B_{q_{3}(t)}\neq 0$ and $Q_{q_{3}(t)}>0$. By taking $|a|<|a^{\prime}_{0}|$, $|x|<|x^{\prime}_{0}|$ small enough, the lower bounds on $Q_{x,y}$ and $B_{x,y}$ follow. ∎
@@ 2.3 Choice of $\theta$
In this subsection, we introduce a number $\theta$ that is going to play a fundamental role in the construction of our combinatorics.
@@
$1<\lambda_{\text{min}}^{2\theta}\mu_{\text{min}}^{3}\text{ and }\lambda_{\text{max}}^{3\theta}\mu^{4}_{\text{max}}<1.$ (2.3.1)
@@
$0<\theta_{0}=\frac{4}{3}\frac{\log\mu_{\text{max}}}{\log{\frac{1}{\lambda_{\text{max}}}}}<\theta<\frac{3}{2}\frac{\log\mu_{\text{min}}}{\log{\frac{1}{\lambda_{\text{min}}}}}=\theta_{1}<\frac{1}{2},$ (2.3.2)
where we used $(F3)$, the initial condition $\lambda_{\text{max}}\mu_{\text{max}}^{3}<1$ and (2.1.7).
Moreover, by adjusting the parameter domain, we can choose $\theta$ such that
$\left|\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right|>w>0.$ (2.3.3)